OpenStudy (anonymous):

There are three containers of equal capacity and all are completely filled with acid and water mixtures in different ratios. The ratioof acid to water in the first container is 2:3, in the second container the ratio is 3:7 and in the third container it is 4:11. If the mixture of all the three containers are mixed together, what is the ratio of acid to water in it?

OpenStudy (shubhamsrg):

let concentration of each container be x then in 1st, acid = 2x/5 water= 3x/5 in 2nd, acid = 3x/10 water = 7x/10 in 3rd, acid = 4x/15 water = 11x/15 when all are mixed, you should be able to find the ratio of total acid/total water//